Solve the following equation for q:
\sqrt[4]{7 q - 9} = 1

Answer by stanbon(75887)

Solve the following equation for q:
\sqrt[4]{7 q - 9} = 1
------
2(7q-9) = 1
7q-9 = 1/2
7q = 19/2
q = 19/14
-----------------
Cheers,
Stan H.
----------------